# LimiX: Unleashing Structured-Data Modeling Capability for Generalist Intelligence

### v1

@@ B.2.2 Generalization for Joint Distribution Learning
Definition B.1 (Approximate Tensorization of Entropy (Caputo & Parisi, 2021)).
@@
Assumption B.1 (Regularity Conditions in the Parametric Class (Li et al., 2024c)).
@@
    There exists $\beta\in(0,1)$ such that $\forall\leavevmode\nobreak\ 1\leq k\leq d,\forall\leavevmode\nobreak\ \pi\in\Pi_{k}$ and $\forall\leavevmode\nobreak\ \theta\in\Theta$, $p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})>0$ implies $q_{\theta}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})>\beta$.
@@
    For any $\epsilon>0$, there exists a partition $\text{Par}(\Theta)=\{\Theta_{1},\dots,\Theta_{|\text{Par}(\Theta)|}\}$ of $\Theta$, such that $\forall\leavevmode\nobreak\ 1\leq k\leq d,\forall\leavevmode\nobreak\ \pi\in\Pi_{k}$, $\forall\leavevmode\nobreak\ \Theta_{i}\in\text{Par}(\Theta)$, $\forall\leavevmode\nobreak\ \theta_{1},\theta_{2}\in\Theta_{i}$, and any $(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$,
@@
The proposition is a corollary from Theorem 4 in Li et al. (2024c), which provides an upper bound for $D_{\operatorname{TV}}(q_{\theta}\parallel p)$ in our setting. The remaining gap is an extension of the result to total variation between conditional distributions. We present Li et al. (2024c)'s result as a lemma.
Lemma B.8 (Li et al. (2024c), Theorem 4).
@@
    There exists $\beta\in(0,1)$ such that $\forall\leavevmode\nobreak\ \pi\subset[d]$ and $\forall\leavevmode\nobreak\ \theta\in\Theta$, $p({\mathbf{x}}_{\pi}|{\mathbf{x}}_{-\pi},\pi)>0$ implies $q_{\theta}({\mathbf{x}}_{\pi}|{\mathbf{x}}_{-\pi},\pi)>\beta$.
@@
    For any $\epsilon>0$, there exists a partition $\text{Par}(\Theta)=\{\Theta_{1},\dots,\Theta_{|\text{Par}(\Theta)|}\}$ of $\Theta$, such that $\forall\leavevmode\nobreak\ \pi\subset[d]$, $\forall\leavevmode\nobreak\ \Theta_{i}\in\text{Par}(\Theta)$, $\forall\leavevmode\nobreak\ \theta_{1},\theta_{2}\in\Theta_{i}$, and any $\mathbf{x}$},
@@
Theorem 4 in Li et al. (2024c) is specified for $\hat{\theta}$ as the minimizer of the empirical loss function. In fact, the proof applies uniformly to arbitrarily $\theta\in\Theta$.
@@
The proof is complete by combining Equations 15 and 16. ∎

### v2

@@ B.2.2 Generalization for Joint Distribution Learning
Definition B.1 (Approximate Tensorization of Entropy (caputo2021block)).
@@
Assumption B.1 (Regularity Conditions in the Parametric Class (li2024promises)).
@@
    There exists $\beta\in(0,1)$ such that $\forall\penalty 10000\ 1\leq k\leq d,\forall\penalty 10000\ \pi\in\Pi_{k}$ and $\forall\penalty 10000\ \theta\in\Theta$, $p(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})>0$ implies $q_{\theta}(\mathbf{X}^{\text{te}}_{\pi}|\mathbf{X}^{\text{ct}},\mathbf{X}^{\text{te}}_{-\pi})>\beta$.
@@
    For any $\epsilon>0$, there exists a partition $\text{Par}(\Theta)=\{\Theta_{1},\dots,\Theta_{|\text{Par}(\Theta)|}\}$ of $\Theta$, such that $\forall\penalty 10000\ 1\leq k\leq d,\forall\penalty 10000\ \pi\in\Pi_{k}$, $\forall\penalty 10000\ \Theta_{i}\in\text{Par}(\Theta)$, $\forall\penalty 10000\ \theta_{1},\theta_{2}\in\Theta_{i}$, and any $(\mathbf{X}^{\text{te}},\mathbf{X}^{\text{ct}})$,
@@
The proposition is a corollary from Theorem 4 in li2024promises, which provides an upper bound for $D_{\operatorname{TV}}(q_{\theta}\parallel p)$ in our setting. The remaining gap is an extension of the result to total variation between conditional distributions. We present li2024promises's result as a lemma.
Lemma B.8 (li2024promises, Theorem 4).
@@
    There exists $\beta\in(0,1)$ such that $\forall\penalty 10000\ \pi\subset[d]$ and $\forall\penalty 10000\ \theta\in\Theta$, $p({\mathbf{x}}_{\pi}|{\mathbf{x}}_{-\pi},\pi)>0$ implies $q_{\theta}({\mathbf{x}}_{\pi}|{\mathbf{x}}_{-\pi},\pi)>\beta$.
@@
    For any $\epsilon>0$, there exists a partition $\text{Par}(\Theta)=\{\Theta_{1},\dots,\Theta_{|\text{Par}(\Theta)|}\}$ of $\Theta$, such that $\forall\penalty 10000\ \pi\subset[d]$, $\forall\penalty 10000\ \Theta_{i}\in\text{Par}(\Theta)$, $\forall\penalty 10000\ \theta_{1},\theta_{2}\in\Theta_{i}$, and any $\mathbf{x}$},
@@
Theorem 4 in li2024promises is specified for $\hat{\theta}$ as the minimizer of the empirical loss function. In fact, the proof applies uniformly to arbitrarily $\theta\in\Theta$.
@@
The proof is complete by combining Equations 17 and 18. ∎